OpenStudy (anonymous):

You roll two dice. What is the probability that the sum of the dice is less than 5 and one die shows a 2?

OpenStudy (anonymous):

(2+0),(2+1),(2+2)= 3 cases total cases = 36 prob. = 3/36 prob. = 1/12

OpenStudy (anonymous):

Thank you